# SPECIAL RELATIVITY

### v9

@@ 1 Introduction
It is generally accepted that covariant equations must have covariant solutions. This statement presented in most textbooks and is obviously incorrect. For instance, from the mathematical viewpoint, the Lorentz covariance of electrodynamics equations is insufficient to guarantee the covariant solutions. We only wish to emphasize here the following point. The principle of special relativity, which says that all physical equations must be invariant under Lorentz transformation (i.e. the pseudo-Euclidean geometry of space-time as an axiom of the theory), does not dictate the reciprocal symmetry of nature.
@@
Authors of textbooks are mistaken in their belief that all inertial frames are equivalent. Relativistic (i.e. reciprocal) symmetry is usually identifies with symmetry under Lorentz boosts. At the same time it is common knowledge that Lorentz boosts alone do not make up a group. The product of two Lorentz boosts in different directions is equal to the product of a pure boost and the spatial (Wigner) rotation. From the mathematical viewpoint, when we are dealing with transformations and set of these transformations does not form a group we have no symmetry between inertial frames. The inertial frames are characterized by motion at constant velocity relative to the fixed stars. There is a real difference between accelerated ( with respect to the fixed stars) inertial frame and the inertial frame without accelerational history and this difference ( in the case of the aberration of light phenomena) is closely associated with the Wigner rotation effect.
@@
It should be note that all well known methods to test the special relativity are second order optical (interference) measurements . The cardinal example is given by the Michelson-Morley experiment. The light from a source is split by a suitable apparatus into two beams and the two overlapping beams produce an interference pattern. The two beam originate in the same source. It is not possible to detect the orbital velocity using second order (interference) experiments. Phase is 4D invariant (i.e. it is simply number). It is independent on the chosen inertial frame and on the chosen coordinatization. This is a consequence of the geometry of space-time. In the interference experiments neither the Doppler effect nor the aberration of light exist separately as well defined physical phenomena and only phase is meaningful quantity. This discussion clearly show that the special relativity calculations yields the observed null fringe shift in the earth-based frame. A close examination of all optical second order experiments inside an inertial frame shows that phenomena appear to be independent of the uniform motion relative to the fixed stars. The usual formulation of the theory of relativity deals with formulated above principle of irrelevancy of velocity, but it should be understood only in a limiting sense when we are dealing with optical second order measurements.
@@
There are a number of other second order experiments. They are not connected with light interference phenomena. The question arises whether it is possible to determine experimentally the state of motion of the initial inertial frame in the accelerated frame by means of time dilation effect for moving reference clock. We demonstrate that slowly down of the accelerated (with respect to the fixed stars) physical clock does not depend on the choice of reference system, in which this effect is measured. In other words, the time dilation effect is absolute, but not relative. It is a fundamental property of the pseudo-Euclidean space-time.
@@
Now let us return to metric diagonalization. In textbooks, time dilation is usually derived from the diagonal (Minkowski) metric. After diagonalization, according to textbooks, the coordinates in inertial frames should be related by the Lorentz transformation. At first site the problem is complete symmetrical and the time dilation effect is relative. Something is evidently wrong. In the Chapter 9 we investigate the connection between time dilation effect and the Langevin metric. Perhaps the only way to demonstrate that the time dilation effect is absolute is to take a mathematical view. From a mathematical viewpoint, the argument that time dilation effect is absolute looks something like the following. It is obvious that metric tensor must be a continuous quantity. The Langevin metric is found by matching the accelerated frame and inertial frame metric tensors. For smooth tailoring of the metric tensor, it is sufficient to relate the coordinates and time of the accelerated observer to the coordinates and time of the inertial observer by the Galilean boost. In this case the metric of inertial frame would pass into the Langevin metric of the accelerated frame. After diagonalization, the metric tensor in the accelerated frame must abruptly change from the Langevin value to the Minkowski (diagonal ) value. We must conclude that when we are dealing with Lorentz transformation and this transformation does not preserve continuity of the metric tensor we have no symmetry between inertial frames. Any change of velocity, or any acceleration (with respect to the fixed stars) has an absolute meaning.
@@
In the Chapter 9-12 we will take up in more detail the absoluteness of acceleration; here we only wished to show how naturally the special relativity leads to the concept of absolute velocity. The formalism of relativistic physics is associated with absolute space-time structure. This structure must be associated with initial inertial frame, which incidentally provides an absolute rest frame. There is no relativity of the uniform motions and we can determine the particle absolute velocity in the initial inertial frame. The main difference between the initial inertial frame and the absolute space of Newton is that special relativity taught us to think in terms of unified space-time model. In special relativity the absolute space and the absolute time of Newton are fused into absolute space-time.
@@
The proper time interval may be identified with particle lifetime, atomic transition periods, nuclear half-life. Until now, all test of time dilation with atomic clocks in the earth-based frame have involved circle path geometry. Indeed, the clocks flown in the commercial jets. When the clocks returned to the laboratory they were compared with the reference clock in the same point at rest with respect to the earth. According to the special relativity, in any time dilation experiment which probe the influence of the orbital velocity one must use the linear (one-way) path geometry. These are experiments of decays of rapidly moving unstable particles. However, the modern experimental techniques are not sensitive enough to register the effect of orbital velocity on the average distance traveled by a disintegrated particle.
@@
In contrast, we describe (see Chapter 8) a simple first order optical experiment within the means of most undergraduate departments that illustrates many of the principles of current research in the foundations of special relativity. It directly demonstrates that one could determine the speed of the earth with respect to the sun-based frame by means of aberration of earth-based incoherent source measurements.
@@
The second part of the book (Chapter 13 - 18) deals with accelerator physics. In the Chapter 13 we describe the particle dynamics method in the Lorentz lab frame using lab time $t$ as evolution parameter. The study of relativistic particle motion in a constant magnetic field according to usual accelerator engineering, is intimately connected with the old (Newtonian) kinematics: the Galilean vectorial law of addition of velocities is actually used. A non-covariant approach to relativistic particle dynamics is based on the absolute time coordinatization, but this is actually a hidden coordinatization. The absolute time synchronization convention is self-evident and this is the reason why this subject is not discussed in accelerator physics.
@@
According to textbooks, if no more than one frame is involved, one does not need to know the theory of relativity. In other words, in this standard approach, introducing as only modification to the classical case the relativistic mass, time differ from space. In convention particle tracking a particle trajectory can be seen from the lab frame as result of successive Galileo boosts that track the motion of the accelerated particle. There is a reason to prefer the non-covariant way within the framework of dynamics only. In this approach we have no mixture of positions and time. This (3+1) dimensional non-covariant particle tracking method is simple, self-evident, and adequate to the laboratory reality. It can be demonstrated that there is no principle difficulty with the non-covariant approach in mechanics and electrodynamics. It is perfectly satisfactory. It does not matter which transformation is used to describe the same reality. What matter is that, once fixed, such convention should be applied and kept in a consistent way for both dynamics and electrodynamics. In fact, the light (group) velocity transforms as a particle velocity. We cannot take old Newton’s kinematics for mechanics and Einstein’s kinematics for electrodynamics.
@@
The common mistake made in accelerator physics is connected with the incorrect algorithm for solving the electromagnetic field equations. If one wants to use the usual Maxwell’s equations, only the solution of the dynamics equations in covariant form (i.e. in Lorentz coordinates) gives the correct coupling between the Maxwell’s equations and particle trajectories in the lab frame.
@@
It is interesting to note that, the use of the conventional coupling of Maxwell’s equations and corrected Newton’s equation for the calculation of the radiation from a moving source does not necessarily leads to mistake. For rectilinear motion of the source and the emitted light beam, the non-covariant and covariant approaches produce the same trajectories, and Maxwell’ equations are compatible with the result of conventional particle tracking. This method was incorrectly extended to the case of non-collinear geometry.
@@
First, we examine the reasoning presented in textbooks. It is generally believed that the magnetic field, in the Lorentz coordinatization, is only capable altering the direction of motion, but not the speed of the electron. Nevertheless, there is argument against this commonly accepted derivation of the composition of velocities. Let us to consider composition of two velocities when they are not lying on a straight line. In conventional particle tracking (similar to the Newton mechanics), one composes them geometrically by the method of parallelogram with velocities regarded as the usual vectors.
@@
Only solution of the dynamics equations in covariant form gives the correct coupling between the usual Maxwell’s equations and particle trajectories in the lab frame. There are two difficulties in solving problems concerning covariant particle tracking. First of all, conventional particle tracking in the single inertial frame is actually based on classical Newton mechanics without reference to Lorentz transformations. Another difficulty is the confusion surrounding the combination of non-parallel velocities in special relativity. The standard textbook presentation of the relativistic velocity transformation is based on the hidden assumption that $(x^{\prime},y^{\prime},z^{\prime})$ axes of the moving observer and $(x,y,z)$ axes of the lab observer are parallel. In other words, it is assumed that observers have common 3-space. This is misconception. The fact that two observers with different trajectories in the Lorentz coordinatization have different 3-spaces is not considered in textbooks. First of all, there is no absolute notion of simultaneity in the theory of special relativity. In the case of the relativity of simultaneity we have a mixture - of positions and time. In other words, in the space measurement of one observer there is mixing of space and time, as seen by the other. Therefore, there is no notion of an instantaneous 3-space.
@@
In textbooks on special relativity, the relativistic composition of velocities is studied analytically from the composition of Lorentz boosts. Authors of textbooks believe that the relativity principle requires that velocity addition law gives the composition law of a group. The non-associativity of the composition law of the non-colinear velocities is not widely known, although it is an exercise in elementary algebra to prove non-associativity. It is therefore impossible to get covariant trajectory, if we have nothing but Lorentz boosts. Something else has to be added. Lorentz boosts do not form a group due to presence of the Wigner rotation. It is comes out quite natural that breakdown of commutativity and associativity in Einstein addition, turn out to be Wigner rotation.
@@
We emphasize that non-collinear velocities addition in the Lorentz coordinatization is regulated by the Wigner rotation. One of the consequences of non-commutativity of non-collinear Lorentz boosts is not the existence of a common ordinary space. Our result is at odds with the prediction from textbooks. The commonly accepted derivation of the composition of perpendicular velocities based on the use the relation $dt^{\prime}=dt/\gamma$ and does not account for the Wigner rotation (i.e. mixture of transverse positions and time). Only the solution of the dynamics equations in covariant form (accounting for the Wigner rotation in the transformation of non-collinear velocities) gives the correct coupling between the Maxwell’s equations and particles trajectories in the lab frame. We illustrate error in standard coupling fields and particles by considering the relatively simple example, wherein the essential physical features are not obscured by unnecessary mathematical difficulties. We hope that the present example will draw wide attention to the central role that Wigner rotation plays in the transformation of non-collinear velocities.
@@
In the Chapter 13 we examine what parts of the dynamics theory depend on the choice of synchronization convention and what parts do not. The present analysis shows the difference between the notions of path and trajectory. The path is rather a purely geometrical notion. The trajectory of a particle conveys more information about its motion because every position is described additionally by the corresponding time instant. The path has exact objective meaning. In fact, the curvature radius of the path in the magnetic field (and consequently the three-momentum) has obviously an objective meaning, i.e. is convention-invariant. In contrast to this, and consistently with the conventionality intrinsic in the velocity, the trajectory $\vec{x}(t)$ of the particle is convention dependent and has no exact objective meaning. Dynamics theory contains a particle trajectory that we do not need to check directly, but which is used in the analysis of electrodynamics problem.
@@
In the Chapter 14 we present a critical reexamination of radiation theory. To evaluate radiation fields arising from an external sources, we need to know the particle velocity $\vec{v}$ and the position $\vec{x}$ as a function of the lab frame time $t$. As discussed above, one should solve the usual Maxwell’s equations in the lab frame with current and charge density created by particles moving along covariant trajectory. For an arbitrary parameter $v/c$ covariant calculations of the radiation process is very difficult. There are, however, circumstances in which calculations can be greatly simplified. As example of such circumstance is a non-relativistic radiation setup. The reason is that the non-relativistic assumption implies the dipole approximation which is of great practical significance. In accounting only for the dipole part of the radiation we neglect all information about the electron trajectory. That means that the dipole radiation does not show any sensitivity to the difference between covariant and non-covariant particle trajectories. But that is only the first and most practically important term. The calculation of higher order corrections to the dipole radiation approximation requires detailed information about the electron trajectory. Obviously, in order to calculate the correction to the dipole radiation, we will have to use the covariant trajectory and not be satisfied with the non-covariant approach.
@@
Accelerator physicists, who try to understand the situation related to the use of the theory of relativity in the synchrotron radiation phenomena, are often troubled by the fact that the difference between covariant and non-covariant particle trajectories was never understood, and that nobody realized that there was a contribution to the synchrotron radiation from relativistic kinematics effects. Accelerator physics was always thought in terms of the old (Newtonian) kinematics that is not compatible with Maxwell’s equations. At this point, a reasonable question arises: since storage rings are designed without accounting for the relativistic kinematics effects, how can they actually operate? In fact, electron dynamics in storage ring is greatly influenced by the emission of radiation. We answer this question in great detail in the Chapter 15. The main emphasis of this chapter is on spontaneous synchrotron radiation from bending magnets and undulators.
@@
Similar to the non-relativistic asymptote, the ultrarelativistic asymptote also provides the essential simplicity of the covariant calculation. The reason is that the ultrarelativistic assumption implies the paraxial approximation. Note that the geometry of the electron motion in a bending magnet has a cylindrical symmetry. There are a number of remarkable effects which are a consequence of the cylindrical symmetry and the paraxial approximation. We demonstrate that there is no needs to use covariant particle tracking for derivation of the bending magnet radiation. However, there is one situation where the conventional theory fails. The covariant approach predicts a non-zero red shift of the critical frequency, which arises when there are perturbations of the electron motion along the magnetic field.
@@
One way to demonstrate incompatibility between the standard approach to relativistic electrodynamics, which deals with the usual Maxwell’s equations, and particle trajectories calculated by using non-covariant particle tracking, is to make a direct laboratory test of synchrotron radiation theory. In other words, we are stating that, despite the many measurements done during decades, synchrotron radiation theory is not an experimentally well-confirmed theory. We analyze the potential for exploiting synchrotron radiation sources in order to confirm the predictions of corrected synchrotron radiation theory. Proposed experiments have never been performed at synchrotron facilities.
@@
The theory of relativity shows that the relativity of simultaneity, which is a most fundamental relativistic kinematics effect, is related with extended relativistic objects. But up to 21 st century there were no macroscopic objects possessing relativistic velocities, and there was a general belief that only microscopic particles in experiments can travel at velocities close to that of light. The 2010s saw a rapid development of new laser light sources in the X-ray wavelength range. An X-ray free electron laser (XFEL) is an example where improvements in accelerator technology makes it possible to develop ultrarelativistic macroscopic objects with an internal structure (modulated electron bunches), and the relativistic kinematics plays an essential role in their description.
@@
In Chapter 16 we present a critical reexamination of existing XFEL theory. It is mainly addressed to readers with limiting knowledge of accelerator physics. We study the production of coherent undulator radiation by modulated ultrarelativistic electron beam. Fortunately, the principle of production of coherent undulator radiation does not require specific knowledge of undulator radiation theory presented in the Chapter 15 and can be explained in a very simple way. Relativistic kinematics enters XFEL physics in a most fundamental way through the rotation of the modulation wavefront, which, in ultrarelativistic approximation, is closely associated to the relativity of simultaneity. When the trajectories of particles calculated in the Lorentz reference frame (i.e. an inertial frame where Einstein synchronization procedure is used to assign values to the time coordinate) they must include relativistic kinematics effects such as relativity of simultaneity. In the ultrarelativistic asymptote, the orientation of the modulation wavefront, i.e the orientation of the plane of simultaneity, is always perpendicular to the electron beam velocity when the evolution of the modulated electron beam is treated using Lorentz coordinates.

### v10

@@ 1 Introduction
It is generally accepted that covariant equations must have covariant solutions. This statement is presented in most textbooks and is incorrect. For instance, from the mathematical viewpoint, the Lorentz covariance of electrodynamics equations is insufficient to guarantee the covariant solutions. We only wish to emphasize here the following point. The principle of special relativity, which says that all physical equations must be invariant under Lorentz transformation (i.e. the pseudo-Euclidean geometry of space-time as an axiom of the theory), does not dictate the reciprocal symmetry of nature.
@@
Authors of textbooks are mistaken in their belief that all inertial frames are equivalent. Relativistic (i.e. reciprocal) symmetry is usually identified with symmetry under Lorentz boosts. At the same time it is common knowledge that Lorentz boosts alone do not make up a group. The product of two Lorentz boosts in different directions is equal to the product of a pure boost and the spatial (Wigner) rotation. From the mathematical viewpoint, when we are dealing with transformations and a set of these transformations does not form a group we have no symmetry between inertial frames. The inertial frames are characterized by motion at constant velocity relative to the fixed stars. There is a real difference between accelerated ( with respect to the fixed stars) inertial frame and the inertial frame without accelerational history and this difference ( in the case of the aberration of light phenomena) is closely associated with the Wigner rotation effect.
@@
It should be noted that all well-known methods to test special relativity are second-order optical (interference) measurements. The cardinal example is given by the Michelson-Morley experiment. The light from a source is split by a suitable apparatus into two beams and the two overlapping beams produce an interference pattern. The two beams originate from the same source. It is not possible to detect the orbital velocity using second-order (interference) experiments. Phase is 4D invariant (i.e. it is simply a number). It is independent of the chosen inertial frame and the chosen coordinatization. This is a consequence of the geometry of space-time. In the interference experiments neither the Doppler effect nor the aberration of light exist separately as well-defined physical phenomena and only phase is a meaningful quantity. This discussion clearly shows that the special relativity calculations yield the observed null fringe shift in the earth-based frame. A close examination of all optical second-order experiments inside an inertial frame shows that phenomena appear to be independent of the uniform motion relative to the fixed stars. The usual formulation of the theory of relativity deals with formulated above principle of irrelevancy of velocity, but it should be understood only in a limiting sense when we are dealing with optical second-order measurements.
@@
There are a number of other second-order experiments. They are not connected with light interference phenomena. The question arises whether it is possible to determine experimentally the state of motion of the initial inertial frame in the accelerated frame using the time dilation effect for a moving reference clock. We demonstrate that the slowdown of the accelerated (with respect to the fixed stars) physical clock does not depend on the choice of reference system, in which this effect is measured. In other words, the time dilation effect is absolute, but not relative. It is a fundamental property of the pseudo-Euclidean space-time.
@@
Now let us return to metric diagonalization. In textbooks, time dilation is usually derived from the diagonal (Minkowski) metric. After diagonalization, according to textbooks, the coordinates in inertial frames should be related to the Lorentz transformation. At first sight, the problem is completely symmetrical and the time dilation effect is relative. Something is evidently wrong. In Chapter 12 we investigate the connection between the time dilation effect and the Langevin metric. Perhaps the only way to demonstrate that the time dilation effect is absolute is to take a mathematical view. From a mathematical viewpoint, the argument that the time dilation effect is absolute looks something like the following. The metric tensor must be a continuous quantity. The Langevin metric is found by matching the accelerated frame and inertial frame metric tensors. For smooth tailoring of the metric tensor, it is sufficient to relate the coordinates and time of the accelerated observer to the coordinates and time of the inertial observer by the Galilean boost. In this case, the metric of the inertial frame would pass into the Langevin metric of the accelerated frame. After diagonalization, the metric tensor in the accelerated frame must abruptly change from the Langevin value to the Minkowski (diagonal ) value. We must conclude that when we are dealing with Lorentz transformation and this transformation does not preserve the continuity of the metric tensor we have no symmetry between inertial frames. Any change of velocity, or any acceleration (with respect to the fixed stars) has an absolute meaning.
@@
The formalism of relativistic physics is associated with absolute space-time structure. This structure must be associated with an initial inertial frame, which incidentally provides an absolute rest frame. There is no relativity of the uniform motions and we can determine the particle’s absolute velocity in the initial inertial frame. The main difference between the initial inertial frame and the absolute space of Newton is that special relativity taught us to think in terms of a unified space-time model. In special relativity, the absolute space and the absolute time of Newton are fused into absolute space-time.
@@
The proper time interval may be identified with particle lifetime, atomic transition periods, and nuclear half-life. Until now, all tests of time dilation with atomic clocks in the earth-based frame have involved circle path geometry. Indeed, the clocks flown in the commercial jets. When the clocks returned to the laboratory they were compared with the reference clock in the same point at rest with respect to the earth. According to special relativity, in any time dilation experiment that probes the influence of the orbital velocity, one must use the linear (one-way) path geometry. These are experiments of decays of rapidly moving unstable particles. However, modern experimental techniques are not sensitive enough to register the effect of orbital velocity on the average distance traveled by a disintegrated particle.
@@
In contrast, we describe (see Chapter 10) a simple first-order optical experiment within the means of most undergraduate departments that illustrates many of the principles of current research in the foundations of special relativity. It directly demonstrates that one could determine the speed of the earth with respect to the sun-based frame using aberration of earth-based incoherent source measurements.
@@
In Chapter 16 we describe the particle dynamics method in the Lorentz lab frame using lab time $t$ as the evolution parameter. The study of relativistic particle motion in a constant magnetic field according to usual accelerator engineering, is intimately connected with the old (Newtonian) kinematics: the Galilean vectorial law of addition of velocities is used. A non-covariant approach to relativistic particle dynamics is based on absolute time coordinatization, but this is a hidden coordinatization. The absolute time synchronization convention is self-evident and this is the reason why this subject is not discussed in accelerator physics.
@@
According to textbooks, if no more than one frame is involved, one does not need to know the theory of relativity. In other words, in this standard approach, introducing only modification to the classical case of the relativistic mass, the time differs from space. In conventional particle tracking a particle trajectory can be seen from the lab frame as a result of successive Galileo boosts that track the motion of the accelerated particle. There is a reason to prefer the non-covariant way within the framework of dynamics only. In this approach, we have no mixture of positions and time. This (3+1) dimensional non-covariant particle tracking method is simple, self-evident, and adequate to the laboratory reality. It can be demonstrated that there is no principal difficulty with the non-covariant approach in mechanics and electrodynamics. It is perfectly satisfactory. It does not matter which transformation is used to describe the same reality. What matters is that once fixed, such a convention should be applied and kept in a consistent way for both dynamics and electrodynamics.
@@
The common mistake made in accelerator physics is connected with the incorrect algorithm for solving the electromagnetic field equations. If one wants to use the Maxwell’s equations, only the solution of the dynamics equations in covariant form (i.e. in Lorentz coordinates) gives the correct coupling between the Maxwell’s equations and particle trajectories in the lab frame.
@@
It is interesting to note that, the use of the conventional coupling of Maxwell’s equations and corrected Newton’s equation for the calculation of the radiation from a moving source does not necessarily lead to a mistake. For the rectilinear motion of the source and the emitted light beam, the non-covariant and covariant approaches produce the same trajectories, and Maxwell’s equations are compatible with the result of conventional particle tracking. This method was incorrectly extended to the case of non-collinear geometry.
@@
First, we examine the reasoning presented in textbooks. It is generally believed that the magnetic field, in the Lorentz coordinatization, is only capable of altering the direction of motion, but not the speed of the electron. Nevertheless, there is an argument against this commonly accepted derivation of the composition of velocities. Let us consider the composition of two velocities when they are not lying on a straight line. In conventional particle tracking (similar to Newton’s mechanics), one composes them geometrically by the method of a parallelogram with velocities regarded as the usual vectors.
@@
Only the solution of the dynamics equations in covariant form gives the correct coupling between the Maxwell’s equations and particle trajectories in the lab frame. There are two difficulties in solving problems concerning covariant particle tracking. First of all, conventional particle tracking in the single inertial frame is based on classical Newton mechanics without reference to Lorentz transformations. Another difficulty is the confusion surrounding the combination of non-parallel velocities in special relativity. The standard textbook presentation of the relativistic velocity transformation is based on the hidden assumption that $(x^{\prime},y^{\prime},z^{\prime})$ axes of the moving observer and $(x,y,z)$ axes of the lab observer are parallel. In other words, it is assumed that observers have a common 3-space. This is a misconception. The fact that two observers with different trajectories in the Lorentz coordinatization have different 3-spaces is not considered in textbooks. First of all, there is no absolute notion of simultaneity in the theory of special relativity. In the case of the relativity of simultaneity, we have a mixture - of positions and time. In other words, in the space measurement of one observer, there is the mixing of space and time, as seen by the other. Therefore, there is no notion of an instantaneous 3-space.
@@
In textbooks on special relativity, the relativistic composition of velocities is studied analytically from the composition of Lorentz boosts. Authors of textbooks believe that the relativity principle requires that the velocity addition law gives the composition law of a group. The non-associativity of the composition law of the non-colinear velocities is not widely known, although it is an exercise in elementary algebra to prove non-associativity. It is therefore impossible to get a covariant trajectory if we have nothing but Lorentz boosts. Something else has to be added. Lorentz boosts do not form a group due to the presence of the Wigner rotation. It comes out quite natural that the breakdown of commutativity and associativity in Einstein’s addition, turns out to be Wigner rotation.
@@
We emphasize that non-collinear velocities addition in the Lorentz coordinatization is regulated by the Wigner rotation. One of the consequences of non-commutativity of non-collinear Lorentz boosts is not the existence of a common ordinary space. Our result is at odds with the prediction from textbooks. The commonly accepted derivation of the composition of perpendicular velocities is based on the use of the relation $dt^{\prime}=dt/\gamma$ and does not account for the Wigner rotation (i.e. mixture of transverse positions and time). Only the solution of the dynamics equations in covariant form (accounting for the Wigner rotation in the transformation of non-collinear velocities) gives the correct coupling between Maxwell’s equations and particle trajectories in the lab frame.
@@
In Chapter 16 we examine what parts of the dynamics theory depend on the choice of synchronization convention and what parts do not. The present analysis shows the difference between the notions of path and trajectory. The path is rather a purely geometrical notion. The trajectory of a particle conveys more information about its motion because every position is described additionally by the corresponding time instant. The path has an exact objective meaning. The curvature radius of the path in the magnetic field (and consequently the three-momentum) has an objective meaning, i.e. is convention-invariant. In contrast to this, and consistently with the conventionality intrinsic in the velocity, the trajectory $\vec{x}(t)$ of the particle is convention-dependent and has no exact objective meaning. Dynamics theory contains a particle trajectory that we do not need to check directly, but which is used in the analysis of electrodynamics problems.
@@
In Chapter 18 we present a critical reexamination of radiation theory. To evaluate radiation fields arising from external sources, we need to know the particle velocity $\vec{v}$ and the position $\vec{x}$ as a function of the lab frame time $t$. As discussed above, one should solve the Maxwell’s equations in the lab frame with current and charge density created by particles moving along the covariant trajectory. For an arbitrary parameter $v/c$ covariant calculations of the radiation process is very difficult. There are, however, circumstances in which calculations can be greatly simplified. An example of such a circumstance is a non-relativistic radiation setup. The reason is that the non-relativistic assumption implies the dipole approximation which is of great practical significance. In accounting only for the dipole part of the radiation we neglect all information about the electron trajectory. That means that the dipole radiation does not show any sensitivity to the difference between covariant and non-covariant particle trajectories. But that is only the first and most practically important term. The calculation of higher-order corrections to the dipole radiation approximation requires detailed information about the electron trajectory. In order to calculate the correction to the dipole radiation, we will have to use the covariant trajectory and not be satisfied with the non-covariant approach.
It is generally assumed that the Maxwell’s equations and corrected Newton’s second law can explain all experiments that are performed in a single inertial (lab) frame. According to textbooks, if no more than one frame is involved, one does not need to know the theory of relativity. Many physicists have already received knowledge about special relativity from textbooks and who would say, ”For those who want to learn just enough about special relativity so they can solve problems - it just changes Newton’s laws by introducing a correction factor to the mass.” This is a misconception. The light velocity transforms as a particle velocity. We cannot take old kinematics for mechanics and Einstein’s kinematics for electrodynamics. We illustrate errors in standard coupling fields and particles by considering the relatively simple example, wherein the essential physical features are not obscured by unnecessary mathematical difficulties. We hope that the present example will draw wide attention to the central role that Wigner rotation plays in the transformation of non-collinear velocities.
@@
Accelerator physicists, who try to understand the situation related to the use of the theory of relativity in the synchrotron radiation phenomena, are often troubled by the fact that the difference between covariant and non-covariant particle trajectories was never understood and that nobody realized that there was a contribution to the synchrotron radiation from relativistic kinematics effects. Accelerator physics was always thought of in terms of the old (Newtonian) kinematics that is not compatible with Maxwell’s equations. At this point, a reasonable question arises: since storage rings are designed without accounting for the relativistic kinematics effects, how can they actually operate? In fact, electron dynamics in the storage ring is greatly influenced by the emission of radiation. We answer this question in great detail in Chapter 19. The main emphasis of this chapter is on spontaneous synchrotron radiation from bending magnets and undulators.
@@
Similar to the non-relativistic asymptote, the ultrarelativistic asymptote also provides the essential simplicity of the covariant calculation. The reason is that the ultrarelativistic assumption implies the paraxial approximation. Note that the geometry of the electron motion in a bending magnet has a cylindrical symmetry. There are a number of remarkable effects which are a consequence of the cylindrical symmetry and the paraxial approximation. We demonstrate that there is no need to use covariant particle tracking for derivation of the bending magnet radiation. However, there is one situation where the conventional theory fails. The covariant approach predicts a non-zero red shift of the critical frequency, which arises when there are perturbations of the electron motion along the magnetic field.
@@
One way to demonstrate incompatibility between the standard approach to relativistic electrodynamics, which deals with the usual Maxwell’s equations, and particle trajectories calculated by using non-covariant particle tracking, is to make a direct laboratory test of synchrotron radiation theory. In other words, we are stating that, despite the many measurements done over decades, synchrotron radiation theory is not an experimentally well-confirmed theory. We analyze the potential for exploiting synchrotron radiation sources to confirm the predictions of corrected synchrotron radiation theory. Proposed experiments have never been performed at synchrotron facilities.
@@
The theory of relativity shows that the relativity of simultaneity, which is a most fundamental relativistic kinematics effect, is related to extended relativistic objects. But up to the 21 st century, there were no macroscopic objects possessing relativistic velocities, and there was a general belief that only microscopic particles in experiments could travel at velocities close to that of light. The 2010s saw a rapid development of new laser light sources in the X-ray wavelength range. An X-ray free electron laser (XFEL) is an example where improvements in accelerator technology make it possible to develop ultrarelativistic macroscopic objects with an internal structure (modulated electron bunches), and the relativistic kinematics plays an essential role in their description.
@@
In Chapter 20 we present a critical reexamination of existing XFEL theory. It is mainly addressed to readers with limited knowledge of accelerator physics. We study the production of coherent undulator radiation by the modulated ultrarelativistic electron beam. Fortunately, the principle of production of coherent undulator radiation does not require specific knowledge of the undulator radiation theory presented in Chapter 19 and can be explained in a very simple way. Relativistic kinematics enters XFEL physics in a most fundamental way through the rotation of the modulation wavefront, which, in ultrarelativistic approximation, is closely associated with the relativity of simultaneity. When the trajectories of particles are calculated in the Lorentz reference frame (i.e. an inertial frame where the Einstein synchronization procedure is used to assign values to the time coordinate) they must include relativistic kinematics effects such as relativity of simultaneity. In the ultrarelativistic asymptote, the orientation of the modulation wavefront, i.e. the orientation of the plane of simultaneity, is always perpendicular to the electron beam velocity when the evolution of the modulated electron beam is treated using Lorentz coordinates.